# Probing supersolidity through excitations in a spin-orbit-coupled Bose-Einstein condensate

### v1

@@ Abstract
Spin-orbit-coupled Bose-Einstein condensates are a flexible experimental platform to engineer synthetic quantum many-body systems. In particular, they host the so-called stripe phase, an instance of a supersolid state of matter. The peculiar excitation spectrum of the stripe phase, a definite footprint of its supersolidity, has so far remained out of experimental reach. We achieve in situ imaging of the stripes and directly observe both superfluid and crystal excitations. We investigate superfluid hydrodynamics and reveal a stripe compression mode, thus demonstrating that the system possesses a compressible crystalline structure. Through the frequency softening of this mode, we locate the supersolid transition point. Our results establish spin-orbit-coupled supersolids as a platform of choice to investigate supersolidity and its rich dynamics.
@@
Supersolidity is an exotic phase of matter characterized by two spontaneously-broken continuous symmetries, translational symmetry and U(1) gauge symmetry [1, 2]. They are associated to two independent order parameters and manifest themselves as a modulated density, reminiscent of a solid, and a global phase coherence, characteristic of a superfluid [3, 4, 5]. This dual nature leads to a remarkably rich excitation spectrum with both superfluid and crystal excitations [6], including two gapless Goldstone modes. After the first realizations of supersolids in quantum gases [7, 8, 9, 10, 11], exploring their dynamics has thus been the focus of experimental research in recent years [12, 13, 14, 15, 16, 17, 18, 19] as it provides an unambiguous signature of supersolidity. Bose-Einstein condensates (BECs) with spin-orbit coupling, where the internal states of the particles are linked to their momentum through optical coupling [20], have been shown to host a supersolid state of matter in the form of the so-called stripe phase. Due to its conceptual simplicity, it can be described analytically and is thus ideally suited to study the exotic properties of supersolids [21, 22, 20, 23, 24]. However, only very fragile realizations of the stripe phase have been demonstrated so far, and the contrast of the stripes, i.e., of the density modulation, was vanishingly small [8, 25]. As a result, the excitations of this phase have been widely believed to be experimentally inaccessible and a debate has arisen over their nature [26]. It was fueled by the observation that other platforms fall into two distinct categories: supersolids that support phonon modes [13, 15], and supersolids that are stiff [12], meaning that the periodicity of their density modulation is fixed. This led to the following question: are spin-orbit-coupled supersolid stripes stiff or do they host phonon modes?
In this work, we experimentally explore the excitation spectrum of the stripe phase and show that spin-orbit-coupled supersolids have a compressible crystal structure and are not stiff. By harnessing the tunable interaction properties of ultracold potassium, we overcome the limitations of previous approaches and realize a stable stripe phase that persists at coupling regimes with strong modulation contrast. While the modulated density pattern could only be probed indirectly in previous experiments [8, 25], here we achieve in situ imaging of the stripes using a matter-wave magnification scheme. By exciting the collective modes of the system, we are able to address both superfluid and crystal-like aspects of supersolids. We show the existence of particle flow between the stripes and observe a stripe compression mode, i.e., an oscillation of the modulation period, which unambiguously proves the compressibility of the crystalline structure. Furthermore, we locate the phase transition from the supersolid to an unmodulated superfluid through frequency softening of the compression mode.
@@ I.   Supersolid spin-orbit-coupled condensates
In spin-orbit-coupled BECs, supersolidity stems from the unconventional dispersion relation and the momentum-dependent interactions that arise due to the mixing of spin and spatial degrees of freedom [21, 20, 22, 23]. The dispersion relation displays two minima in momentum space [28], which results in two possible phases at the many-body level, sketched in Fig. 1(a). The plane-wave phase, which appears when a single minimum is occupied, has a homogeneous density profile. In contrast, the stripe phase, which occurs when both minima are occupied, has a spontaneously modulated density profile while it retains the spatial phase coherence of the condensate [21, 22, 20, 23], and constitutes a supersolid [6]. The experimental challenge is to create a situation where the stripe phase occupies a significant portion of the phase diagram [20]. To quantitatively investigate the stability of the stripe phase and find a suitable regime to probe supersolid excitations, we have developed an analytical model.
This model considers the dressed states at the two minima of the dispersion relation, which emerges when two different states $\downarrow$ and $\uparrow$ (e.g., internal atomic states [20, 25], or different bands of an optical lattice [8]) are coupled through an optical field that transfers momentum to the atoms. A pair of laser beams in a two-photon Raman configuration with frequency difference $\Delta\omega_{\mathrm{R}}$, two-photon Rabi frequency $\Omega$ and two-photon detuning $\delta$ is used, as depicted in Fig. 1(b). It leads to a momentum transfer $2\hbar k_{\mathrm{R}}$ along the $x$ direction $\mathbf{e_{x}}$, with $\hbar k_{\mathrm{R}}$ the recoil momentum unit and $\hbar=h/(2\pi)$ the reduced Planck’s constant. The atom-photon dressed states have a momentum-dependent spin composition, which is set by the polarization parameter $P=\tilde{\delta}/\tilde{\Omega}$, where $\tilde{\delta}=\delta-2\hbar k_{\mathrm{R}}k/m$ and $\tilde{\Omega}=\sqrt{\Omega^{2}+\tilde{\delta}^{2}}$ are the generalized detuning and Rabi frequency. Here $m$ is the atomic mass and $\hbar k$ is the momentum along $x$ in the frame rotating at $\Delta\omega_{\mathrm{R}}$ and co-moving with the Raman coupling fields [29]. In this frame, the single-particle Hamiltonian of the system takes the translationally invariant form
where $\sigma_{x}$ and $\sigma_{z}$ are Pauli matrices. For $\hbar\Omega<4\,E_{\mathrm{R}}$, where $E_{\mathrm{R}}=\hbar^{2}k_{\mathrm{R}}^{2}/2m$ is the Raman recoil energy, the dispersion relation of the lower dressed state features two minima, see Fig. 1(c). The system behaves as a mixture of condensates in the left ($\ell$) and right ($r$) wells that are non-orthogonal. Thus, a modulated density profile emerges due to interference between them when both $\ell$ and $r$ are simultaneously occupied and miscible [28, 20]. Their overlap increases with the spin-orbit coupling strength, leading to a contrast of the density modulation that, away from the transition, scales linearly with $\Omega$ [29].
Figure 2: In situ observation of supersolid stripes. (a) Single shot experimental image of the in situ density profile of the cloud at $\hbar\Omega=2.00\pm 0.08\,E_{\mathrm{R}}$ measured with matter-wave optics. The image is rescaled to represent the expected cloud aspect ratio before matter-wave magnification along $x$. Top right: corresponding density profile integrated along the stripes. The black line denotes a modulated Gaussian fit with stripe spacing $d$. Bottom right: histogram showing the spatial phase of the modulation $\phi$ over 350 realizations. (b) Spacing $d$ of the stripes vs. Raman coupling $\Omega$ measured from in situ (blue circles) and time-of-flight (gray squares, see [29]). Solid line: single-particle theory prediction, obtained from the inverse of the momentum difference of the minima of the dispersion relation (shown on the right for different spin-orbit coupling strengths $\Omega$). Errorbars correspond to one standard error of the mean (eom). The blue shaded area shows the systematic error of the in situ data from the magnification calibration.
Our effective field theory is based on the effective interactions of the optically-dressed states, which are given by the polarization at the dispersion minima [29]. We name it the mixture model because it builds on the mixture picture introduced by [28, 20]. It delivers analytic expressions for both ground state properties and collective excitations [29] and highlights the conceptual simplicity of spin-orbit-coupled supersolids. Figure 1(d) shows the effective interaction parameters according to our mixture model for a gas with bare scattering lengths $a_{\uparrow\uparrow}$, $a_{\downarrow\downarrow}$, and $a_{\uparrow\downarrow}$ at detuning $\delta=0$. Increasing $\Omega$ reduces the effective intra-well interactions $a_{\ell\ell}$, $a_{rr}$ and increases the inter-well ones $a_{\ell r}$. As a result, the miscibility of the system is reduced and it becomes energetically favorable to occupy a single minimum of the dispersion relation as soon as $a_{\ell r}>a_{\ell\ell}$ or $a_{\ell r}>a_{rr}$. The crystal-like modulation is then lost and the system transitions to the plane-wave state. This relation between interactions and coupling strength demonstrates that obtaining a stripe phase at large couplings requires $a_{\uparrow\downarrow}\ll a_{\uparrow\uparrow},a_{\downarrow\downarrow}$. This situation is not available in atomic species like the 87Rb used in previous works, where the small scattering length difference yields a very fragile supersolid phase with a contrast of the density modulation $C<5\%$ [20, 30, 31, 25]. However, we can achieve it by exploiting two-component potassium BECs, where the interactions are tunable using Feshbach resonances [32]. The complete phase diagram of the system with 41K at 51.7 G in the parameter space of $\Omega$ and $\delta$ is depicted in Fig. 1(e). The stripe phase appears as a dome that is not centered at $\delta=0$ due to the setting $a_{\downarrow\downarrow}<a_{\uparrow\uparrow}$. The phase diagram shows the contrast $C$ according to the variational model of Refs. [23, 27], in excellent agreement with the black dashed lines, which are analytical predictions from our mixture model. This constitutes an unusually large, stable, and high-contrast supersolid stripe phase and provides an excellent regime to perform our experiments.
II.   In situ observation of supersolid stripes
We realize the supersolid stripe phase with a Bose-Einstein condensate of 41K atoms in states ||$\ket{\downarrow}\equiv\ket{F=1,m_{F}=0}$ and ||$\ket{\uparrow}\equiv\ket{F=1,m_{F}=-1}$. By setting the magnetic field to $B=51.7\pm 0.1$ G, between an intra-spin and an inter-spin Feshbach resonance, we tune the bare state interactions to $(a_{\uparrow\uparrow},a_{\downarrow\downarrow},a_{\uparrow\downarrow})=(115\pm 10,65\pm 1,40\pm 8)\,a_{0}$, where $a_{0}$ is the Bohr radius [32]. We couple the two states with two Raman laser beams of orthogonal polarizations at the tune-out wavelength $\lambda_{\mathrm{R}}=768.97$ nm, for which scalar potentials cancel. We work at a zero two-photon Raman detuning $\hbar\delta=0.00\pm 0.02\,E_{\mathrm{R}}$ (gray band in Fig. 1(e)), where the phase transition point is independent of the atomic density for our parameter regime [33]. The Raman beams have a relative angle of $\theta=94\pm 1^{\circ}$, which leads to a momentum transfer along the $x$ direction of $2\hbar k_{\mathrm{R}}=4\pi\hbar\mathrm{sin}(\theta/2)/\lambda_{\mathrm{R}}$ [29].
@@ II.   In situ observation of supersolid stripes
To investigate the periodicity of the stripes, we extract the stripe spacing from in situ images taken at different spin-orbit coupling strengths $\Omega$. Figure 2(b) shows that the stripe spacing $d$ increases significantly with $\Omega$, which can be understood from a momentum-space picture. The stripe spacing is given by the momentum difference $\Delta k$ of the two dressed states through $d=2\pi/\Delta k$ which, at equilibrium, corresponds to the difference between the two minima of the dispersion relation (see right panel). It can also be inferred from spin-resolved momentum-space images obtained through time-of-flight expansion. Contrary to the in situ one, this analysis is model dependent [29], resulting in a spacing $\pi/k_{\mathrm{R}}$ at the zero-coupling point by construction. The data is shown as gray squares in Fig. 2(b). Within their respective uncertainties, both in situ and time-of-flight data agree with the theory prediction. Thus, our observations demonstrate that the period of spin-orbit-coupled supersolids is not externally imposed by the wavelength of the coupling laser. This variable crystalline pattern stands in contrast to cavity-mediated supersolidity, where the modulation period is fixed by the light wavelength [12].
@@ III.   Probing superfluid hydrodynamics through excitations
Figure 3: Crystal Goldstone mode and superfluid hydrodynamics. (a) Population imbalance between side stripes $\eta$ vs. displacement of the central stripe $D$ with respect to the center of mass of the cloud (see inset and [29]). For a harmonically-trapped system with superfluid flow, where a sliding of the density modulation is compensated by particle exchange between the stripes, both quantities are linearly correlated and provide evidence for the crystal Goldstone mode. (b) Top panel: Breathing mode of the overall cloud, observed through its size $\sigma$ vs. time. Solid line: sinusoidal fit used to extract the breathing frequency $\omega_{\mathrm{B}}$, yielding $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=1.5\pm 0.1\sim\sqrt{5/2}$ with $\omega_{\mathrm{D}}$ the dipole mode frequency, as expected from superfluid hydrodynamics. Bottom panel: number of stripes in the supersolid cloud. It dynamically changes during half of a breathing oscillation (see shaded area in top panel), supporting the existence of superfluid flow across the system. Errorbars denote one standard eom.
@@
Supersolids are characterized by a superfluid flow through the crystalline structure that maintains the global phase coherence of the system [5, 1]. While previous experiments demonstrated phase coherence in spin-orbit-coupled supersolids using Talbot interferometry [25], our in situ images allow us to investigate instead the superfluid hydrodynamic properties of the system. To this end, we followed the procedure developed for supersolids in dipolar gases [14] and studied the population imbalance between stripes. Figure 3(a) shows the imbalance between side stripes $\eta$ as a function of the displacement $D$ of the central stripe with respect to the center of mass of the cloud (see illustration of $D$ and $\eta$ in the inset and [29]). We see a clear linear correlation between displacement and imbalance, revealing that a phase shift of the crystal structure is compensated by a superfluid counterflow. Such a phase shift provides evidence that the low-energy crystal Goldstone mode of the finite-size system is excited [14].
@@
To probe other signatures of superfluidity, we investigate the collective breathing and dipole modes of the system, which we detect through the evolution of the cloud’s width $\sigma$ and the displacement of its center of mass $x_{\mathrm{0}}$, respectively. In conventional BECs, a ratio of their frequencies $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=\sqrt{5/2}$ (with $\omega_{\mathrm{B}}$ and $\omega_{\mathrm{D}}$ the breathing and dipole mode frequencies) signals a system described by superfluid hydrodynamics and thus supports the superfluid character of the cloud [34]. Fig. 3(b) shows the breathing oscillation of our system. We find a frequency of $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=1.5\pm 0.1$, in good agreement with the conventional BEC prediction. This behavior is consistent with numerical simulations of the Gross-Pitaevskii equation which implicitly assume superfluid hydrodynamics [33]. Here the dipole mode frequency $\omega_{\mathrm{D}}$ differs from the trapping frequency $\omega_{x}$ because spin-orbit-coupled systems violate Kohn’s theorem [35, 36], analogously to superfluids subjected to an optical lattice [37, 38]. Another interesting aspect is that, due to the changing size of the cloud, the number of stripes in the supersolid varies with time. The bottom panel of Fig. 3(b) shows that the superfluid flow can dynamically create additional stripes. Thus, our data support the superfluid character of the system via the frequency of its breathing mode and the existence of dynamic particle flow between the stripes.

### v2

@@ Abstract
Spin-orbit-coupled Bose-Einstein condensates are a flexible experimental platform to engineer synthetic quantum many-body systems. In particular, they host the so-called stripe phase, an instance of a supersolid state of matter. The peculiar excitation spectrum of the stripe phase, a definite footprint of its supersolidity, has been difficult to measure experimentally. Here, we perform in situ imaging of the stripes and directly observe both superfluid and crystal excitations. We investigate superfluid hydrodynamics and reveal a stripe compression mode, thus demonstrating that the system possesses a compressible crystalline structure. Through the frequency softening of this mode, we locate the supersolid transition point. Our results establish spin-orbit-coupled supersolids as ideal systems to investigate supersolidity and its rich dynamics.
@@
Supersolidity is an exotic phase of matter characterized by two spontaneously-broken continuous symmetries, translational symmetry and U(1) gauge symmetry [1, 2]. They are associated with two independent order parameters and manifest themselves, respectively, as a modulated density, reminiscent of the crystalline structure of a solid, and a global phase coherence, characteristic of a superfluid [3, 4, 5]. In recent years, several quantum-gas platforms have realized superfluids with a spontaneous density modulation [6, 7, 8, 9, 10]. Among these, Bose-Einstein condensates (BECs) with spin-orbit coupling, where the internal states of the particles are linked to their momentum through optical coupling [11], have been shown to host a supersolid state of matter where the density modulation takes the form of stripes. Owing to its conceptual simplicity, this so-called stripe phase is analytically tractable and serves as a powerful platform for exploring supersolid phenomena [12, 13, 11, 14, 15].
Figure 1: A robust high-contrast spin-orbit-coupled supersolid. (a) Spin-orbit-coupled BECs at small coupling have a dispersion relation with two minima, leading to two possible many-body phases: the unmodulated plane-wave phase (one minimum occupied) and the modulated supersolid-stripe phase (both minima occupied). (b) Two-photon Raman coupling scheme with two beams of frequency difference $\Delta\omega_{\mathrm{R}}$, coupling strength $\Omega$, detuning $\delta$, and recoil momentum per beam $\hbar k_{\mathrm{R}}$ (tunable through the angle of the Raman beams $\theta$), used to engineer spin-orbit coupling. (c) Dispersion relation $E(k)$ of the lower optically-dressed state (solid line), compared to that of the bare atomic states (dashed lines). Colorscale: spin polarization $P$ of the dressed state. The system can be modeled as a mixture of condensates in the left $\ell$ and right $r$ wells, which consist of a coherent superposition of the bare spin states (red and blue circles) and have modified effective interaction properties given by the effective scattering lengths $a_{\ell\ell}$, $a_{\ell r}$ and $a_{rr}$. (d) Effective interaction parameters at $\delta=0$ vs. Raman coupling $\Omega$, calculated via the mixture model. Vertical black dashed line: supersolid phase transition point $\Omega_{\mathrm{c}}$ separating the supersolid-stripe phase and the plane-wave phase. This panel shows how tuning interaction strengths can stabilize the stripe phase vs. the plane-wave phase. (e) Phase diagram of a homogeneous spin-orbit-coupled BEC featuring a large and stable supersolid stripe phase, calculated with the variational ansatz of [14, 16]. $C$ is the contrast of the density modulation. Dashed lines: phase boundaries calculated with the mixture model. Gray band: region experimentally explored in this work, with detuning $\hbar\delta=0.00\pm 0.02E_{\mathrm{R}}$. Parameters in (d) and (e) correspond to 41K at a magnetic field $B=51.7$ G and with a density $1\text{\times}{10}^{14}\text{\,}\text{atoms}\mathrm{/}\mathrm{c}\mathrm{m}^{3}$, which realizes a robust and high-contrast supersolid stripe phase.
In supersolids, the existence of two order parameters leads to two gapless Goldstone modes and to a plethora of superfluid and crystal excitations [17]. Probing this remarkably rich excitation spectrum has become a central topic of research across the different quantum-gas systems displaying supersolid properties [18, 19, 20, 21, 22, 23, 24, 25]. Recently, the compressibility of the supersolids' emergent crystal structure has been put forward as an additional requirement for full-fledged supersolidity, given that solids naturally host crystal phonons. This property has been observed in the supersolid phase of dipolar quantum gases [19, 21, 23], whereas it is lacking in BECs where effective long-range interactions are induced by single-mode cavities [6, 18]. Addressing this question in spin-orbit-coupled BECs has been challenging. Because of the extreme fragility of the realized stripe states, characterized by a vanishingly small stripe contrast [7, 26], the excitations of the supersolid phase in these systems have been considered experimentally inaccessible, and its properties have been widely debated [27]. By utilizing spin-orbit-coupled potassium Bose-Einstein condensates, where the relative strengths of interspin and intraspin interactions are tunable through Feshbach resonances [28], we overcome the limitations of previous approaches and realize a stable stripe phase that persists at coupling regimes with strong modulation contrast. We are thus able to experimentally explore its excitation spectrum, show that its crystal structure is compressible, and precisely locate the supersolid phase transition.
@@ I. Supersolid spin-orbit-coupled condensates
In spin-orbit-coupled BECs, supersolidity stems from the unconventional dispersion relation and the momentum-dependent interactions that arise thanks to the mixing of spin and spatial degrees of freedom [12, 11, 13, 14]. The dispersion relation displays two minima in momentum space [29], which results in two possible phases at the many-body level (Fig. 1(a)). The plane-wave phase, which appears when a single minimum is occupied, has a homogeneous density profile. In contrast, the stripe phase, which occurs when both minima are occupied, has a spontaneously modulated density profile while it retains the spatial phase coherence of the condensate [12, 13, 11, 14], which constitutes a supersolid [17]. The experimental challenge is to create a situation where the stripe phase occupies a significant portion of the phase diagram [11]. To quantitatively investigate the stability of the stripe phase and find a suitable regime to probe supersolid excitations, we have developed an analytical model.
This model considers the dressed states at the two minima of the dispersion relation, which emerges when two different states $\downarrow$ and $\uparrow$ (e.g., internal atomic states [11, 26], or different bands of an optical lattice [7]) are coupled through an optical field that transfers momentum to the atoms. To realize this coupling, we use a pair of laser beams in a two-photon Raman configuration with frequency difference $\Delta\omega_{\mathrm{R}}$, two-photon Rabi frequency $\Omega$ and two-photon detuning $\delta$ (Fig. 1(b)). It leads to a momentum transfer $2\hbar k_{\mathrm{R}}$ along the $x$ direction $\mathbf{e_{x}}$, with $\hbar k_{\mathrm{R}}$ the recoil momentum unit and $\hbar=h/(2\pi)$ the reduced Planck constant. The atom-photon dressed states have a momentum-dependent spin composition, which is set by the polarization parameter $P=\tilde{\delta}/\tilde{\Omega}$, where $\tilde{\delta}=\delta-2\hbar k_{\mathrm{R}}k/m$ and $\tilde{\Omega}=\sqrt{\Omega^{2}+\tilde{\delta}^{2}}$ are the generalized detuning and Rabi frequency. Here $m$ is the atomic mass and $\hbar k$ is the momentum along $x$ in the frame rotating at $\Delta\omega_{\mathrm{R}}$ and co-moving with the Raman coupling fields [30]. In this frame, the single-particle Hamiltonian of the system takes the translationally invariant form
where $\sigma_{x}$ and $\sigma_{z}$ are the Pauli matrices. For $\hbar\Omega<4\,E_{\mathrm{R}}$, where $E_{\mathrm{R}}=\hbar^{2}k_{\mathrm{R}}^{2}/2m$ is the Raman recoil energy, the dispersion relation of the lower dressed state features two minima (Fig. 1(c)). The system behaves as a mixture of condensates in the left ($\ell$) and right ($r$) wells that are non-orthogonal. Thus, a modulated density profile emerges owing to an interference between them when both $\ell$ and $r$ are simultaneously occupied and miscible [29, 11]. As the spin-orbit coupling strength increases, $\ell$ and $r$ become less orthogonal to each other. This larger wavefunction overlap leads to an increased contrast of the density modulation which, away from the transition, scales linearly with $\Omega$ [30].
Figure 2: In situ observation of supersolid stripes. (a) Single shot experimental image of the in situ density profile of the cloud at $\hbar\Omega=2.00\pm 0.08\,E_{\mathrm{R}}$ measured with matter-wave optics. The image is rescaled to represent the expected cloud aspect ratio before matter-wave optics magnification along $x$. Top right: corresponding density profile integrated along the stripes. The black line denotes a modulated Gaussian fit with stripe spacing $d$. Bottom right: histogram showing the spatial phase of the modulation $\phi$ over 350 realizations. (b) Spacing $d$ of the stripes vs. Raman coupling $\Omega$ measured from in situ (blue circles) and time-of-flight (gray squares, see [30]). Solid line: single-particle theory prediction, obtained from the inverse of the momentum difference of the minima of the dispersion relation (shown in the inset for different spin-orbit coupling strengths $\Omega$) [30]. Errorbars correspond to one standard error of the mean (eom). The blue shaded area shows the systematic error of the in situ data from the magnification calibration.
Our effective field theory is based on the effective interactions of the optically-dressed states, which are given by the polarization at the dispersion minima [30]. We name it the mixture model because it builds on the mixture picture introduced in [29, 11]. It delivers analytical expressions for both ground state properties and collective excitations [30] and highlights the conceptual simplicity of spin-orbit-coupled supersolids. Figure 1(d) shows the effective interaction parameters according to our mixture model for a gas with bare scattering lengths $_{\uparrow\uparrow}, aa_{\downarrow\downarrow}$, and $a_{\uparrow\downarrow}$ at detuning $\delta=0$. Increasing $\Omega$ reduces the effective intra-well interactions $a_{\ell\ell}$, $a_{rr}$ and increases the inter-well ones $a_{\ell r}$. As a result, the miscibility of the system is reduced and it becomes energetically favorable to occupy a single minimum of the dispersion relation as soon as $a_{\ell r}>a_{\ell\ell}$ or $a_{\ell r}>a_{rr}$. The crystal-like modulation is then lost and the system transitions to the plane-wave state. This relation between interactions and coupling strength demonstrates that obtaining a stripe phase at large couplings requires $a_{\uparrow\downarrow}\ll a_{\uparrow\uparrow},a_{\downarrow\downarrow}$. This situation is not available in atomic species like the 87Rb used in previous works, where the small scattering length difference yields a very fragile supersolid phase with a contrast of the density modulation $C<5\%$ [11, 31, 32, 26]. However, we can achieve it by exploiting two-component potassium BECs, where the interspin interactions are tunable using a Feshbach resonance [28]. The complete phase diagram of the system with 41K at 51.7 G in the parameter space of $\Omega$ and $\delta$ is depicted in Fig. 1(e). The stripe phase appears as a dome that is not centered at $\delta=0$ owing to the setting $a_{\downarrow\downarrow}<a_{\uparrow\uparrow}$. The phase diagram shows the contrast $C$ according to the variational model of Refs. [14, 16], in excellent agreement with the black dashed lines, which are analytical predictions from our mixture model. This constitutes an unusually large, stable, and high-contrast supersolid stripe phase and provides an excellent regime in which to perform our experiments.
II. In situ observation of supersolid stripes
@@
To investigate the periodicity of the stripes, we extract the stripe spacing from in situ images taken at different spin-orbit coupling strengths $\Omega$. Figure 2(b) shows that the stripe spacing $d$ increases significantly with $\Omega$, which can be understood from a momentum-space picture. The stripe spacing is given by the momentum difference $\Delta k$ of the two dressed states through $d=2\pi/\Delta k$ which, at equilibrium, corresponds to the difference between the two minima of the dispersion relation (Fig. 2(b), inset). It can also be inferred from spin-resolved momentum-space images obtained through time-of-flight expansion (Fig. S4) [30]. Contrary to the in situ analysis, the analysis of the time-of-flight data is model-dependent, resulting in a spacing $\pi/k_{\mathrm{R}}$ at the zero-coupling point by construction. The data are shown as gray squares in Fig. 2(b). Within their respective uncertainties, both in situ and time-of-flight data agree with the theory prediction. Thus, our observations demonstrate that the period of spin-orbit-coupled supersolids is not externally imposed by the wavelength of the coupling laser. This variable crystalline pattern stands in contrast to that of cavity systems, where the modulation period is fixed by the light wavelength [18].
@@ III. Probing superfluid hydrodynamics through excitations
Supersolids are characterized by a superfluid flow through the crystalline structure that maintains the global phase coherence of the system [5, 1]. Whereas previous experiments demonstrated phase coherence in spin-orbit-coupled supersolids using Talbot interferometry [26], our in situ images allow us to investigate instead the superfluid hydrodynamic properties of the system. To this end, we followed the procedure developed for supersolids in dipolar gases [20] and studied the population imbalance between stripes. Figure 3(a) shows the imbalance between side stripes $\eta$ as a function of the displacement $D$ of the central stripe with respect to the center of mass of the cloud (see illustration of $D$ and $\eta$ in the inset and [30]). We see a clear linear correlation between displacement and imbalance, revealing that a phase shift of the crystal structure is compensated by a superfluid counterflow. Such a phase shift provides evidence that the low-energy crystal Goldstone mode of the finite-size system is excited [20].
Figure 3: Crystal Goldstone mode and superfluid hydrodynamics. (a) Population imbalance between side stripes $\eta$ vs. displacement of the central stripe $D$ with respect to the center of mass of the cloud (see inset and [30]). For a harmonically-trapped system with superfluid flow, where a sliding of the density modulation is compensated by particle exchange between the stripes, the two quantities are linearly correlated and provide evidence for the crystal Goldstone mode. (b) Top panel: breathing mode of the overall cloud, observed through its size $\sigma$ vs. time. Solid line: sinusoidal fit used to extract the breathing frequency $\omega_{\mathrm{B}}$, yielding $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=1.5\pm 0.1\sim\sqrt{5/2}$ with $\omega_{\mathrm{D}}$ the dipole mode frequency, as expected from superfluid hydrodynamics. Bottom panel: number of stripes in the supersolid cloud. It dynamically changes during half of a breathing oscillation (see shaded area in top panel), supporting the existence of superfluid flow across the system. Errorbars denote one standard eom.
@@
To probe other signatures of superfluidity, we investigate the collective breathing and dipole modes of the system, which we detect through the evolution of the cloud's width $\sigma$ and the displacement of its center of mass $x_{\mathrm{0}}$, respectively. In conventional BECs, a ratio of their frequencies $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=\sqrt{5/2}$ (with $\omega_{\mathrm{B}}$ and $\omega_{\mathrm{D}}$ the breathing and dipole mode frequencies) signals a system described by superfluid hydrodynamics and thus supports the superfluid character of the cloud [34]. Figure 3(b) shows the breathing oscillation of our system. We find a frequency of $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=1.5\pm 0.1$, in good agreement with the conventional BEC prediction. This behavior is consistent with numerical simulations of the Gross-Pitaevskii equation which implicitly assume superfluid hydrodynamics [33]. Here the dipole mode frequency $\omega_{\mathrm{D}}$ differs from the trapping frequency $\omega_{x}$ because spin-orbit-coupled systems violate Kohn's theorem [35, 36], analogously to superfluids subjected to an optical lattice [37, 38]. Another notable aspect is that, owing to the changing size of the cloud, the number of stripes in the supersolid varies with time. The bottom panel of Fig. 3(b) shows that the superfluid flow can dynamically create additional stripes. Thus, our data support the superfluid character of the system via the frequency of its breathing mode and the existence of dynamic particle flow between the stripes.